A questionnaire asks shareholders of a company to state whether they consider the chairman’s salary to be too high, about right, or too low. Excluding shareholders who have no opinions, the probabilities of answers from random selected shareholders are as follows:

Too high 0.82

About right 0.12

Too low. 0.05


1
Expert's answer
2021-10-18T13:34:41-0400

What are the probabilities that if three shareholders are selected at random, that they will all answer, too high?


"P=(P(Too\ high))^3=0.82^3=0.5514"